OpenStudy (anonymous):

If l = 12 cm and m = 5 cm, what is the length of n?

OpenStudy (anonymous):

diagram?

OpenStudy (anonymous):

helpp

OpenStudy (austinl):

Pythagorean Theorem, \(a^2+b^2=c^2\Rightarrow l^2+m^2=n^2\)

OpenStudy (anonymous):

^yes just put the values into the Pythagorean Theorem, \(a^2+b^2=c^2\), which in this case is \(l^2+m^2=n^2\) \[(12)^2+(5)^2=n^2\]\[144+25=n^2\]\[169=n^2\]\[\sqrt{169}=\sqrt{n^2}\]\[13=n\]and so , the third side, \(n\), is \(13cm\)